# Criminal organizations exhibit hysteresis, resilience, and robustness by balancing security and efficiency

### v2

@@ Abstract
00footnotetext: Corresponding author: c.vanelteren@uva.nl22footnotetext: Institute for Advanced Study, University of Amsterdam, Amsterdam, The Netherlands33footnotetext: Polder center, Institute for Advanced Study, The Netherlands44footnotetext: Computational science lab, University of Amsterdam, The Netherlands
@@
This study explores the underlying mechanisms that govern criminal network formation, expansion, and potential dissolution. Using a model based on concepts from evolutionary game theory, we analyze these networks as collaborative assemblies of roles, considering expected costs, potential benefits, and the certainty of expected outcomes. Our experiments investigate how the interplay of dynamics on social networks and changes to environmental conditions impact the robustness of criminal organizations against law enforcement efforts. The approach complements the extensive literature on multi-actor coordination games by applying them to a criminal context [undefy, undefz, undefaa, undefab, undefac, undefad, undefae, undefaf].
@@ 1.1 Strategy update
where $\epsilon$ represents the decision error controlling the agent’s uncertainty when changing strategy. Notice that the transition or conditional probability of changing to a given strategy is independent of the base state, mapping directly to a multinomial logit model that reflects a law of relative effect [undefag] and, for a single role, maps to the classic Glauber dynamics for the kinetic Ising models [undefah].
@@ 2.1 Dynamics of efficient criminal organizations
fig.˜1 shows the fixed points of the dynamics of efficient criminal organizations, which characterize where the system is attracted to (stable fixed points) or repelled from (unstable fixed points). In the case of high decision error, high $\epsilon$, the system exhibits a single stable fixed point, which attracts the dynamics to an intermediate fraction of criminals. In this regime with a single equilibrium, increasing the cost-to-benefit ratio consistently and reversibly decreases the fraction of criminals. When the decision increases further, all dynamics resort to a random choice regardless of the cost of the criminal act, e.g., compare $\epsilon=\infty$ to $\epsilon=5$. This reflects the conditions in which the lines between criminal and non-criminal acts are blurred. An example could be the formation of a "second" government such as the prevalence of the Mafia in southern Italy [undefai]. However, reducing the decision error can either decrease the fraction of criminals if the cost-to-benefit ratio is high enough or increase it for a low cost-to-benefit ratio. The number of roles, $R$, reflecting the complexity of the task, controls the critical cost-to-benefit ratio that determines the direction of the effect (precisely, $(c/b)^{\mathrm{crit}}=1/2^{R-1}$). For low decision error, low $\epsilon$, the dynamics may exhibit three fixed points: two stable attractors, one with a high fraction of criminals (e.g., the points between A and B) and one with low (e.g., the points between C and D), and an unstable state at an intermediate fraction of criminals. For a high cost-to-benefit ratio, above $(c/b)^{\mathrm{H}}$, only the attractor with a low fraction of criminals is present; for low cost-to-benefit ratio, below $(c/b)^{\mathrm{L}}$, only the attractor with a high fraction of criminals is present; and for intermediate values, between $(c/b)^{\mathrm{L}}$ and $(c/b)^{\mathrm{R}}$, the three are present and the dynamics can be attracted to two different stable states. When the system is in a stable state with a high fraction of criminals (A), changes to the cost-to-benefit ratio are largely ineffective in changing the fraction of criminals ($A\to B$) unless they surpass $(c/b)^{\mathrm{R}}$ ($B\to C$). Similarly, when the system is in a stable state with a low fraction of criminals (C), a reduction of the cost-to-benefit will not lead to an increase ($C\to D$) unless the reduction is beyond $(c/b)^{\mathrm{L}}$ ($D\to C$). In the regime where the three states are present, with intermediate values of cost-to-benefit ratio, the initial number of criminals determines where the system is attracted. The unstable fixed point determines the critical mass of criminals below which criminal activity collapses to the low criminal state and above which it grows to the high criminal state. As the cost-to-benefit ratio increases, this critical mass also increases.
@@ 2.2 State-dependent fortitude: hysteresis in criminal organizations
Firstly, we consider fluctuations in the fraction of criminals, focusing again on fig.˜1. Over time, the number of criminals may change due to various factors, such as deaths, rivalries, and other dynamics. The ability to resist fluctuations in the population will be denoted as resilience [undefaj, undefm].
Secondly, the viability of a criminal market is influenced by the relative market value of the product being sold and different risks. Fluctuations in supply and demand and law enforcement surveillance impact the sustainability of criminal organizations. The ability to resist fluctuations due to external pressure is denoted as robustness[undefaj, undefm]. For this, we will use fig.˜2.
@@ 3 Discussion & conclusions
Network analysis provides an intuitive approach to decompose the structure of a criminal organization in terms of security and efficiency. However, the results of this article highlight the need to move beyond mere structural analysis and consider the intricate relationship between the organizational structure and the emergence, as well as the stability of criminal organizations amidst external conditions and interventions. Exploring the interplay between cost-benefit analysis and decision-making errors provides insights into how criminal organizations maintain resilience and robustness amidst external perturbations. These findings complement existing empirical research in criminology by providing underlying mechanisms driving the proliferation of illicit enterprises and informing strategic approaches for combating organized crime [undefp, undefo, undefak, undeff, undefg, undefh, undefal].
@@
The relationship between stricter punishment and its impact on organized crime is complex and subject to debate among experts [undefam, undefan, undefao, undefal]. While deterrence theory suggests that stricter punishment can reduce criminal behavior [undefap], its effectiveness in deterring organized crime is not always clear-cut due to the sophisticated strategies employed by criminal organizations, such as bribery, corruption, and intimidation [undefaq]. Moreover, the lack of effectiveness of increased punishment on organized crime may be attributed to the creation of in-group versus out-group dynamics that occur within criminal organizations. In such circumstances criminal co-offence reduces the impact of external perturbations and increases their resilience and robustness [undefi, undefar].
Increased punishment is not the sole determinant of the effectiveness of a criminal organization [undefas]. Depending on the diversity of the portfolio of the criminal organization, the benefits derived from engaging in criminal activities are contingent upon the forces of supply and demand [undefd]. These forces may compel criminal organizations to alter their criminal portfolio, either by moving into different financial markets or by changing their methods of operation [undefat, undefau].
The insights gained from this study have important implications for policymakers and law enforcement agencies. They can be used to inform the development of targeted intervention strategies that address network resilience, awareness dynamics, and societal factors driving illicit behavior. The results reveal how interventions of similar scale may lead to the downfall of a criminal organization under certain conditions, while delivering marginal effects under others. This implies that increasing criminal punishment may not have the desired effects on reducing organized crime, while having a profound effect on the personal freedoms novel policies may entail [undefav, undefaw, undefax]. Understanding the cost-benefit analysis and decision-making processes within criminal organizations is imperative for formulating effective policy frameworks aimed at preventing and disrupting organized crime [undefay, undefaz].
Furthermore, this perspective reframes law enforcement policies not merely as reactive measures to criminal actions but as integral components of a dynamic system. Within this system, policymakers, law enforcement agencies, and criminal organizations form a collective wherein each party can influence the other. This systemic approach underscores the interplay between enforcement strategies and criminal behavior, emphasizing the need for proactive, adaptive policy measures that anticipate and counteract criminal activities effectively [undefaaa, undefaab, undefp, undefo].
@@ A.1 Graph generation and link density, and assortativity
where $M$ is the joint probability distribution or mixing matrix [undefaac].
@@ A.4 Criminal opportunity
see [undefaad] for proof.
The model formulation is akin to a n-order Ising model with base $\{0,1\}$ under transversal magnetic field where each agent takes local averages [undefaae].
@@ Appendix E Theories of organized crime
I shall not today attempt further to define the kinds of material I understand to be embraced within that shorthand description. But I know it when I see it, and the motion picture involved in this case is not that I know it when I see it - Justice Potter Stewart [undefaaf]
@@ E.2 Illicit enterprise theory
The value chain includes all roles or activities that support the primary activities bringing a raw resource or service to market [undefaag], human resource management, research and development, firm infrastructure and so on. The value chain describes the business process or coarse steps that make up the business.
@@ E.3 Social embededness of criminal networks
For a given representation of a social network, other subnetworks can be identified. Consider, for example, a network formed from contact data. Within this social network different subnetworks can be identified. The network of work colleagues is embedded on this social network, similar to friend networks, or kinship networks. The concept of social embedding gains its origins from social economy [undefaah]. Sociologists identified that economic transactions are embedded within the networks of personal relations. Criminal networks, therefore, are another form of embedding [undeff].
The general idea is that criminal opportunity is formed through social relations. Social relations do not happen at random but often abey the laws of social an geographic distance [undefaai]. The closer people live, the more daily activity people have in common, the less social distance between them. This results in a clustering of people based on factors such as geographic distance, ethnicity, education, age, membership of a local sports club and so on. This aspect is also found in criminal networks [undefaaj].

### v3

@@ Abstract
00footnotetext: Corresponding author: c.vanelteren@uva.nl11footnotetext: Computational Science Lab, Informatics Institute, University of Amsterdam, The Netherlands22footnotetext: Institute for Advanced Study, University of Amsterdam, Amsterdam, The Netherlands33footnotetext: POLDER Center, Institute for Advanced Study, The Netherlands
@@
This study explores the underlying mechanisms that govern criminal network formation, expansion, and potential dissolution. Using a model based on concepts from evolutionary game theory, we analyze these networks as collaborative assemblies of roles, considering expected costs, potential benefits, and the certainty of expected outcomes. Our experiments investigate how the interplay of dynamics on social networks and changes to environmental conditions impact the robustness of criminal organizations against law enforcement efforts. The approach complements the extensive literature on multi-actor coordination games by applying them to a criminal context [undefy, undefz, undefaa, undefab, undefac, undefad, undefae, undefaf, undefag, undefah].
@@ 1.1 Strategy update
where $\epsilon$ represents the decision error controlling the agent’s uncertainty when changing strategy. Notice that the transition or conditional probability of changing to a given strategy is independent of the base state, mapping directly to a multinomial logit model that reflects a law of relative effect [undefai] and, for a single role, maps to the classic Glauber dynamics for the kinetic Ising models [undefaj].
@@ 2.1 Dynamics of efficient criminal organizations
fig.˜1 shows the fixed points of the dynamics of efficient criminal organizations, which characterize where the system is attracted to (stable fixed points) or repelled from (unstable fixed points). In the case of high decision error, high $\epsilon$, the system exhibits a single stable fixed point, which attracts the dynamics to an intermediate fraction of criminals. In this regime with a single equilibrium, increasing the cost-to-benefit ratio consistently and reversibly decreases the fraction of criminals. When the decision increases further, all dynamics resort to a random choice regardless of the cost of the criminal act, e.g., compare $\epsilon=\infty$ to $\epsilon=5$. This reflects the conditions in which the lines between criminal and non-criminal acts are blurred. An example could be the formation of a "second" government such as the prevalence of the Mafia in southern Italy [undefak]. However, reducing the decision error can either decrease the fraction of criminals if the cost-to-benefit ratio is high enough or increase it for a low cost-to-benefit ratio. The number of roles, $R$, reflecting the complexity of the task, controls the critical cost-to-benefit ratio that determines the direction of the effect (precisely, $(c/b)^{\mathrm{crit}}=1/2^{R-1}$). For low decision error, low $\epsilon$, the dynamics may exhibit three fixed points: two stable attractors, one with a high fraction of criminals (e.g., the points between A and B) and one with low (e.g., the points between C and D), and an unstable state at an intermediate fraction of criminals. For a high cost-to-benefit ratio, above $(c/b)^{\mathrm{H}}$, only the attractor with a low fraction of criminals is present; for low cost-to-benefit ratio, below $(c/b)^{\mathrm{L}}$, only the attractor with a high fraction of criminals is present; and for intermediate values, between $(c/b)^{\mathrm{L}}$ and $(c/b)^{\mathrm{R}}$, the three are present and the dynamics can be attracted to two different stable states. When the system is in a stable state with a high fraction of criminals (A), changes to the cost-to-benefit ratio are largely ineffective in changing the fraction of criminals ($A\to B$) unless they surpass $(c/b)^{\mathrm{R}}$ ($B\to C$). Similarly, when the system is in a stable state with a low fraction of criminals (C), a reduction of the cost-to-benefit will not lead to an increase ($C\to D$) unless the reduction is beyond $(c/b)^{\mathrm{L}}$ ($D\to C$). In the regime where the three states are present, with intermediate values of cost-to-benefit ratio, the initial number of criminals determines where the system is attracted. The unstable fixed point determines the critical mass of criminals below which criminal activity collapses to the low criminal state and above which it grows to the high criminal state. As the cost-to-benefit ratio increases, this critical mass also increases.
@@ 2.2 State-dependent fortitude: hysteresis in criminal organizations
Firstly, we consider fluctuations in the fraction of criminals, focusing again on fig.˜1. Over time, the number of criminals may change due to various factors, such as deaths, rivalries, and other dynamics. The ability to resist fluctuations in the population will be denoted as resilience [undefal, undefm].
Secondly, the viability of a criminal market is influenced by the relative market value of the product being sold and different risks. Fluctuations in supply and demand and law enforcement surveillance impact the sustainability of criminal organizations. The ability to resist fluctuations due to external pressure is denoted as robustness[undefal, undefm]. For this, we will use fig.˜2.
@@ 3 Discussion & conclusions
Network analysis provides an intuitive approach to decompose the structure of a criminal organization in terms of security and efficiency. However, the results of this article highlight the need to move beyond mere structural analysis and consider the intricate relationship between the organizational structure and the emergence, as well as the stability of criminal organizations amidst external conditions and interventions. Exploring the interplay between cost-benefit analysis and decision-making errors provides insights into how criminal organizations maintain resilience and robustness amidst external perturbations. These findings complement existing empirical research in criminology by providing underlying mechanisms driving the proliferation of illicit enterprises and informing strategic approaches for combating organized crime [undefp, undefo, undefam, undeff, undefg, undefh, undefan].
@@
The relationship between stricter punishment and its impact on organized crime is complex and subject to debate among experts [undefao, undefap, undefaq, undefan]. While deterrence theory suggests that stricter punishment can reduce criminal behavior [undefar], its effectiveness in deterring organized crime is not always clear-cut due to the sophisticated strategies employed by criminal organizations, such as bribery, corruption, and intimidation [undefas]. Moreover, the lack of effectiveness of increased punishment on organized crime may be attributed to the creation of in-group versus out-group dynamics that occur within criminal organizations. In such circumstances criminal co-offence reduces the impact of external perturbations and increases their resilience and robustness [undefi, undefat].
Increased punishment is not the sole determinant of the effectiveness of a criminal organization [undefau]. Depending on the diversity of the portfolio of the criminal organization, the benefits derived from engaging in criminal activities are contingent upon the forces of supply and demand [undefd]. These forces may compel criminal organizations to alter their criminal portfolio, either by moving into different financial markets or by changing their methods of operation [undefav, undefaw].
The insights gained from this study have important implications for policymakers and law enforcement agencies. They can be used to inform the development of targeted intervention strategies that address network resilience, awareness dynamics, and societal factors driving illicit behavior. The results reveal how interventions of similar scale may lead to the downfall of a criminal organization under certain conditions, while delivering marginal effects under others. This implies that increasing criminal punishment may not have the desired effects on reducing organized crime, while having a profound effect on the personal freedoms novel policies may entail [undefax, undefay, undefaz]. Understanding the cost-benefit analysis and decision-making processes within criminal organizations is imperative for formulating effective policy frameworks aimed at preventing and disrupting organized crime [undefaaa, undefaab].
Furthermore, this perspective reframes law enforcement policies not merely as reactive measures to criminal actions but as integral components of a dynamic system. Within this system, policymakers, law enforcement agencies, and criminal organizations form a collective wherein each party can influence the other. This systemic approach underscores the interplay between enforcement strategies and criminal behavior, emphasizing the need for proactive, adaptive policy measures that anticipate and counteract criminal activities effectively [undefaac, undefaad, undefp, undefo].
@@ A.1 Graph generation and link density, and assortativity
where $M$ is the joint probability distribution or mixing matrix [undefaae].
@@ A.4 Criminal opportunity
see [undefaaf] for proof.
The model formulation is akin to a n-order Ising model with base $\{0,1\}$ under transversal magnetic field where each agent takes local averages [undefaag].
@@ Appendix E Theories of organized crime
I shall not today attempt further to define the kinds of material I understand to be embraced within that shorthand description. But I know it when I see it, and the motion picture involved in this case is not that I know it when I see it - Justice Potter Stewart [undefaah]
@@ E.2 Illicit enterprise theory
The value chain includes all roles or activities that support the primary activities bringing a raw resource or service to market [undefaai], human resource management, research and development, firm infrastructure and so on. The value chain describes the business process or coarse steps that make up the business.
@@ E.3 Social embededness of criminal networks
For a given representation of a social network, other subnetworks can be identified. Consider, for example, a network formed from contact data. Within this social network different subnetworks can be identified. The network of work colleagues is embedded on this social network, similar to friend networks, or kinship networks. The concept of social embedding gains its origins from social economy [undefaaj]. Sociologists identified that economic transactions are embedded within the networks of personal relations. Criminal networks, therefore, are another form of embedding [undeff].
The general idea is that criminal opportunity is formed through social relations. Social relations do not happen at random but often abey the laws of social an geographic distance [undefaak]. The closer people live, the more daily activity people have in common, the less social distance between them. This results in a clustering of people based on factors such as geographic distance, ethnicity, education, age, membership of a local sports club and so on. This aspect is also found in criminal networks [undefaal].